A 10 W resistor ,15.9 mH inductor and 159 micro farad capacitor are connected in parallel to a 200V, 50 Hz source. Calculate the supply current.


1
Expert's answer
2021-05-31T17:02:10-0400

Gives 

Resistance (R)=10 ohm

Power supply (V)=200V

Frequency (n)=50Hz

Inductor(L)=15.9mH

Capacitor (C)=159microferad

"I=\frac{V}{Z}"

Z=impedence

"Z=\sqrt{R^2+({X_L-X_C})^2}"

XL=wL="2\pi f" "X_L=2\times3.14\times50\times15.9\times10^{-3}=4.9926 \Omega"


"X_C=\frac{1}{2\times3.14\times50\times159\times10^{-6}}=20.03\Omega"

put Value

"Z=\sqrt{10^2+15.03^2}=18.05\Omega"

"I=\frac{200}{18.05}=11.08A"